# Robust Policy Switching for Antifragile Reinforcement Learning for UAV Deconfliction in Adversarial Environments

### v1

@@ Abstract
The increasing automation of navigation for unmanned aerial vehicles (UAVs) has exposed them to adversarial attacks that exploit vulnerabilities in reinforcement learning (RL) through sensor manipulation. Although existing robust RL methods aim to mitigate such threats, their effectiveness has limited generalization to out-of-distribution shifts from the optimal value distribution, as they are primarily designed to handle fixed perturbation. To address this limitation, this paper introduces an antifragile RL framework that enhances adaptability to broader distributional shifts by incorporating a switching mechanism based on discounted Thompson sampling (DTS). This mechanism dynamically selects among multiple robust policies to minimize adversarially induced state-action-value distribution shifts. The proposed approach first derives a diverse ensemble of action-robust policies by accounting for a range of perturbations in the policy space. These policies are then modeled as a multiarmed bandit (MAB) problem, where DTS optimally selects policies in response to nonstationary Bernoulli rewards, effectively adapting to evolving adversarial strategies. Theoretical framework has also been provided where by optimizing the DTS to minimize the overall regrets due to distributional shift, results in effective adaptation against unseen adversarial attacks thus inducing antifragility. Extensive numerical simulations validate the effectiveness of the proposed framework in complex navigation environments with multiple dynamic three-dimensional obstacles and with stronger projected gradient descent (PGD) and spoofing attacks. Compared to conventional robust, non-adaptive RL methods, the antifragile approach achieves superior performance, demonstrating shorter navigation path lengths and a higher rate of conflict-free navigation trajectories compared to existing robust RL techniques.
@@ I Introduction
Unmanned aerial vehicles (UAVs) have become pivotal in tasks such as aerial surveillance [huang2022aerial, zheng2019evolutionary], target tracking [zhang2018coarse], reconnaissance [obermeyer2009path], and logistics [dissanayaka2023review]. As their deployment in civil airspace expands, ensuring safe and conflict-free navigation is critical [huang2019collision], necessitating robust unmanned aircraft traffic management (UTM) systems [kopardekar2016unmanned]. According to ICAO [organisation2005global], conflict detection and resolution (CDR) in UTM operates across three layers: strategic conflict management (pre-flight deconfliction) [tang2023incorporating, tsourdos23], separation provision (in-flight tactical decisions) [hamissi2023survey, Huang2023UTM], and collision avoidance (post-conflict mitigation) [yasin2020unmanned]. While strategic deconfliction offers early conflict prevention, it lacks flexibility to handle sudden changes like airspace closures. Tactical deconfliction, by contrast, supports real-time adaptation through onboard sensors and surveillance technologies.
Reinforcement learning (RL) has emerged as a promising tool for tactical deconfliction [zhao2021physics, li2024graph], enabling UAVs to adapt to real-time conditions by learning optimal control policies in high-dimensional state-action spaces. However, RL policies are heavily reliant on sensor data and communication channels, which are vulnerable to adversarial attacks such as spoofing, jamming, and evasion [yu2023cybersecurity, apruzzese2020deep, panda2023fragility, Kai24spoof]. These threats not only degrade perception but can also exploit the RL agent’s policy directly [Kai2024RL, wei2024survey].
To address these challenges, this paper explores an antifragile RL framework for UAV tactical deconfliction under adversarial attacks. Rather than merely withstanding disruptions, the proposed method adapts and improves from them by leveraging an ensemble of robust policies and dynamically selecting among them using distributional feedback. Our approach integrates adversarially trained policies with Discounted Thompson Sampling (DTS) to adaptively switch to the most robust policy in response to unseen perturbations, enhancing both safety and resilience in UTM environments.
@@ I-A1 Tactical UAV Deconfliction
Traditional deconfliction approaches combine trajectory optimization with cost functions that adapt to dynamic environments. Rule-based planners such as A* [guan2023autonomous], Dijkstra’s algorithm [zhang2020trajectory], and artificial potential fields (APF) [sun2017collision] offer limited adaptability in real time. More sophisticated frameworks using model predictive control (MPC) [ji2016path] and mixed-integer linear programming (MILP) [alonso2010collision] improve obstacle avoidance but remain computationally intensive and often limited to 2D environments.
Reinforcement learning (RL) offers a scalable alternative for 3D tactical deconfliction by optimizing policies over high-dimensional state spaces. However, existing RL-based methods often struggle with generalization under adversarial perturbations. To enhance efficiency and safety in 3D navigation, fluid dynamics-based trajectory planning, such as interfered fluid dynamics systems (IFDS) [yao2015uav, yao2015real], has been used to incorporate UAV kinematics and reduce path cost [celestini2022trajectory]. Recent work integrating IFDS with deep RL has shown improved resilience and adaptability [zhang2021adaptive]. Building upon this, our approach enhances RL-based deconfliction by incorporating adaptive policy switching to counter adversarial threats.
@@ I-A2 Adversarial Threats and Robust RL Defenses
Adversarial attacks in RL are commonly categorized into perturbations of the state, action, model, and reward spaces [ilahi2021challenges, Kai2024RL, chu2024survey]. State-space attacks use gradient-based loss functions [pattanaik2018robust] or saliency maps [papernot2016limitations] to mislead perception. Action-space and model-space attacks manipulate policy behavior and model integrity [lee2020spatiotemporally, huai2020malicious], while reward-space attacks distort agent learning [han2018reinforcement, Kai24spoof]. Defensive strategies include adversarial training [pattanaik2018robust, behzadan2017whatever], defensive distillation [papernot2016distillation, delgrange2022distillation], and risk-aware modeling [pan2019risk, tessler2019action]. These methods enhance robustness by preparing agents for known perturbation types. Other techniques include noisy reward shaping [wang2020reinforcement], falsified temporal logic constraints [pinto2017robust], and adversary-aware training [zhang2021robust].
However, most robust RL approaches assume stationary or bounded distribution shifts. When exposed to previously unseen or evolving threats, their performance often degrades due to lack of adaptability. Static defenses cannot dynamically respond to adversarial strategies that change over time. Our work addresses this limitation by inducing adaptability through policy switching, enabling the agent to operate effectively across diverse threat scenarios.
I-A3 Antifragility in Autonomous Systems
Antifragility refers to the ability of a system not only to resist disruption but to improve under it [taleb2013mathematical]. In contrast to robustness, which ensures performance preservation, antifragile systems leverage perturbations as feedback for adaptation. Recent work in robotics and transportation systems explores antifragile design via uncertainty-aware control [sun2024exploring, axenie2023antifragile], fragility spectrum quantification [pravin2024fragility], and structural resilience in AAM networks [altun2023analyzing]. Despite these advances, most approaches isolate antifragility from robustness rather than modeling them as points on a resilience continuum [de2020antifragility]. Additionally, the lack of a balancing feedback mechanism limits practical antifragility. In our approach, this feedback is achieved by switching among robust models trained under varying perturbation conditions, allowing the agent to improve its response under previously unseen attacks while preserving core safety objectives.
I-B Novelty, Contribution and Impact
In this paper, we present a novel antifragility framework for UAV deconfliction in environments consisting of 3D obstacles, integrating adaptive switching between pre-trained adversarial action-robust RL policies and a standard pre-trained policy. First, we design action-robust RL policies that account for various perturbations in the policy space. To dynamically select the most appropriate policy in response to adversarial threats, we employ a discounted Thompson sampling (DTS) approach, modeling robust policies as multi-armed bandits (MAB). The switching mechanism is designed to minimize the value distribution shifts caused by adversarial attacks, which are generated using the Frank-Wolfe algorithm. The DTS-based switching strategy enables the system to select robust policies effectively in the presence of previously unseen adversarial attacks. The key contributions of this work are as follows.
  • Antifragile Policy Switching Framework: We propose a novel antifragile reinforcement learning framework for UAV deconfliction in 3D obstacle environments, combining an ensemble of action-robust policies with a Discounted Thompson Sampling (DTS)-based switching strategy. Robust policies are modeled as arms in a multi-armed bandit (MAB) setup, and selected adaptively based on value distributional shifts induced by adversarial attacks.
  • Theoretical Guarantees for Robust Adaptation: We present formal regret analysis and convergence guarantees showing that DTS asymptotically selects the optimal robust policy under nonstationary, unseen adversarial perturbations. Additionally, we theoretically characterize antifragility as the agent’s ability to improve performance under increasing distributional shifts via adaptive policy switching.
  • Empirical Validation in Adversarial UAV Scenarios: We benchmark the proposed approach against existing robust RL methods, including adversarial and probabilistic training, under projected gradient descent (PGD) and hijack-induced spoofing attacks in complex 3D airspace environments. Our method consistently outperforms baselines in reward preservation and conflict avoidance.
@@ I-B Novelty, Contribution and Impact
Paper Organization:Section II elaborates on the RL and IFDS methodologies used for deconfliction in environments with 3D obstacles. Section E.II provides a comprehensive examination of action-based RL. In Section IV, the formulation of the robust policies as multi-armed bandits are considered, along with the respective Bernouli rewards. Section V focuses on the optimality of the Bernoulli rewards with respect to the distribution shift considering the regret minimization problem, and the optimality of adaptation during test-time due to DTS sampler while inducing antifragility. A detailed analysis of the results, including a comparative assessment against robust algorithms based on benchmarks, is provided in Section VI. Finally, Section VII synthesizes the key findings and conclusions.
Figure 1: The navigation environment for UAV against 3D dynamic obstacle with the trajectory computed using IFDS.
@@ II-A State Transition Model
The modelling environment is a tactical deconfliction problem as shown in Fig.1, where the UAV has to navigate from the starting point to the goal via a dynamic 3D obstacle. The resulting trajectory is computed by the interaction between the original flow field due to the target and the interfered flow field due to the obstacle(s). The initial flow field is designed to converge at the target destination where position vector $\mathbf{P}$, influences the velocity $u\left(\mathbf{P}\right)$ of the UAV as:
@@
where $C$ is the convergence speed and $\bar{d}$ being the Euclidean distance between current position $\mathbf{P}$ and the target position $\mathbf{P}_{d}$.
@@
However, the initial flow field is affected due to the presence of dynamic obstacles. The extent of the change can be captured using $\overline{M}\left(\mathbf{P}\right)$ which is derived in Appendix A, which is dependant on the shape of the obstacles. The resulting synthetic disturbance flow field in the dynamic obstacle environment is provided as follows,
@@
For the point $\mathbf{P}_{t}$, on the path $\mathbf{P}_{t+1}$ is obtained by integrating the disturbed flow field:
@@
In reality, the IFDS manifests itself in the modified actions after the output of the optimization agent action space (e.g., in the rotorcraft controllers) as per the implementation shown in Figure 2.
@@ II-B Reinforcement Learning Agent
The first step in defining the RL problem involves formulating the environment transition model given by the position update equation in (3), which accounts for both the initial and perturbed flow fields. The subsequent subsections will explain the state, action, and reward for the RL agent.
@@ II-B2 Reward
As described in (6), the deconfliction is achieved by prioritizing the shortest path to the target. However, this shortest path may bring the UAV closer to the obstacle surface. To address this risk, a threat zone is defined around the obstacle, where a penalty is imposed if the planned path intersects this zone.
@@ II-B3 Action
The parameters $\left\{\varrho_{0},\varsigma_{0},\vartheta_{n}\right\}$ for the IFDS as an action of the agent.
@@ III Action Robust Reinforcement Learning
As illustrated in Figure 3, the schematic is shown to obtain a set of robust policies by varying the parameter $\alpha$. The agent and adversarial policies, denoted as $\pi_{\theta}$ and $\mu_{\omega}$, respectively, generate action $a_{t}$ based on a weighted combination: $\left(1-\alpha_{i}\right)a^{\textup{agent}}$ for the agent and $\alpha a^{\textup{adv}}$ for the adversary. These actions determine the trajectory of navigation of a UAV in 3D obstacle environment. Upon executing an action $a_{t}$, the agent receives a reward $r_{t}$ while transitioning to a new state $\Phi^{\prime}$. The agent’s overall experience is stored as $e_{t}=\left\{\Phi_{t},a_{t},r_{t},\Phi^{\prime}_{t}\right\}$ in the replay buffer $\mathcal{D}_{\textup{replay}}$. Subsequently, a batch of experiences $\mathcal{D}=\left\{e_{1},\cdots,e_{N}\right\}$ is sampled from the replay buffer to compute the critic and the corresponding actor loss function.
@@
However, to train agent and adversarial policies in a shared action space, a two-player zero-sum Markov game is used [tessler2019action]. In this setting, both players simultaneously select an action and the reward is determined by the current state $\Phi_{t}$ and the joint actions of the agent and adversarial policies, which influences the transition dynamics. The MDP models the game as $\mathcal{M}_{2}=\left(\mathcal{S},\mathcal{A},\mathcal{A}^{\prime},T_{2},\gamma,R_{2},P_{0}\right)$ where $\mathcal{A}$ and $\mathcal{A}^{\prime}$ denote the action spaces available to the agent and the adversary. The state transition probability is given by $T_{2}:\mathcal{S}\times\mathcal{A}\times\mathcal{A}^{\prime}\times\mathcal{S}\rightarrow\mathbb{R}$ and $R_{2}:\mathcal{S}\times\mathcal{A}\times\mathcal{A}^{\prime}\rightarrow\mathbb{R}$ defines the reward for both players. Here, the agent’s policy is represented by $\mu:\mathcal{S}\rightarrow\mathcal{A}$, while the adversary follows the policy $\nu:\mathcal{S}\times\mathcal{A}^{\prime}$ within the environment $\mathcal{M}_{2}$. At each time step $t$, the agent and the adversary select actions $a_{t}=\mu\left(\Phi_{t}\right)$ and ${a_{t}}^{\prime}=\nu\left(\Phi_{t}\right)$, respectively. In the context of a zero-sum game, the agent receives a reward $r_{t}=R_{2}\left(\phi_{t},a_{t},a^{{}^{\prime}}_{t}\right)$, while the adversary receives a reward of $-r_{t}$. In the adversarial game, the following performance objective is considered for optimization.
@@
[…]

### v3

# Meta Policy Switching for Secure UAV Deconfliction in Adversarial Airspace

@@ Abstract
Autonomous UAV navigation using reinforcement learning (RL) is vulnerable to adversarial attacks that manipulate sensor inputs, potentially leading to unsafe behavior and mission failure. Although robust RL methods provide partial protection, they often struggle to generalize to unseen or out-of-distribution (OOD) attacks due to their reliance on fixed perturbation settings. To address this limitation, we propose a meta-policy switching framework in which a meta-level policy dynamically selects among multiple robust policies to counter unknown adversarial shifts. At the core of this framework lies a discounted Thompson sampling (DTS) mechanism that formulates policy selection as a multi-armed bandit problem, thereby minimizing value distribution shifts via self-induced adversarial observations. We first construct a diverse ensemble of action-robust policies trained under varying perturbation intensities. The DTS-based meta-policy then adaptively selects among these policies online, optimizing resilience against self-induced, piecewise-stationary attacks. Theoretical analysis shows that the DTS mechanism minimizes expected regret, ensuring adaptive robustness to OOD attacks and exhibiting emergent antifragile behavior under uncertainty. Extensive simulations in complex 3D obstacle environments under both white-box (Projected Gradient Descent) and black-box (GPS spoofing) attacks demonstrate significantly improved navigation efficiency and higher conflict-free trajectory rates compared to standard robust and vanilla RL baselines, highlighting the practical security and dependability benefits of the proposed approach.
@@ I Introduction
With the rising deployment of UAVs in civil airspace, ensuring safe and conflict-free navigation is critical [huang2019collision], which necessitates robust unmanned aircraft traffic management (UTM) systems [kopardekar2016unmanned]. According to ICAO [organisation2005global], conflict detection and resolution (CDR) in UTM operates across three layers: strategic conflict management (pre-flight deconfliction) [tsourdos23], separation provision (in-flight tactical decisions) [Huang2023UTM], and collision avoidance (post-conflict mitigation) [yasin2020unmanned]. While strategic deconfliction offers early conflict prevention, it lacks flexibility to handle sudden changes like dynamic obstacles with unknown trajectories. Tactical deconfliction, by contrast, supports real-time adaptation through onboard sensors and surveillance technologies.
Reinforcement learning (RL) has emerged as a promising tool for tactical deconfliction [zhao2021physics, li2024graph], enabling UAVs to adapt to real-time conditions by learning optimal control policies in high-dimensional state-action spaces. However, RL policies are heavily reliant on sensor data and communication channels, which are vulnerable to adversarial attacks such as spoofing, jamming, and evasion [yu2023cybersecurity, panda2023fragility, Kai24spoof]. These threats not only degrade perception but can also exploit the RL agent’s policy directly [Kai2024RL, wei2024survey]. To address these challenges, this paper explores a meta-policy switching framework for UAV tactical deconfliction under adversarial attacks. Rather than merely withstanding disruptions, the meta-policy learns to switch against an ensemble of robust policies against self-induced adversarial observations. Our approach integrates adversarially trained policies with discounted Thompson sampling (DTS) to adaptively switch to the most robust policy in response to unseen perturbations, enhancing both safety and resilience in UTM environments.
@@ I-A1 Tactical UAV Deconfliction
Reinforcement learning (RL) offers a scalable alternative for tactical deconfliction over existing planning algorithms by optimizing policies over high-dimensional state spaces to train obtain navigation policy against dynamic obstacles. However, existing RL-based methods often struggle with generalization under adversarial perturbations. To enhance efficiency and safety against 3D obstacles, techniques such as interfered fluid dynamics systems (IFDS) [yao2015uav, yao2015real], has been used to incorporate UAV kinematics and reduce path cost [celestini2022trajectory]. Recent work integrating IFDS with deep RL has shown improved resilience and adaptability [zhang2021adaptive]. Building upon this, our approach enhances RL-based deconfliction by incorporating adaptive policy switching to counter adversarial threats.
@@ I-A2 Adversarial Threats and Robust RL Defenses
Adversarial attacks in RL are broadly categorized into perturbations of the action and state space, policy network parameters, and reward functions [ilahi2021challenges, Kai2024RL, chu2024survey]. State-space attacks use gradient-based loss functions [pattanaik2018robust] or saliency maps [papernot2016limitations] to mislead the trained policies. Action-space and policy network parameter attacks manipulate policy behavior and model integrity [lee2020spatiotemporally, huai2020malicious], while reward-space attacks distort agent learning [Kai24spoof]. Defensive strategies enhance robustness by preparing the agents against known perturbation bounds which include adversarial training [pattanaik2018robust, behzadan2017whatever], defensive distillation [papernot2016distillation, delgrange2022distillation], risk-aware modeling [pan2019risk, tessler2019action], noisy reward shaping [wang2020reinforcement] and robust adversary-aware RL (RARL) [zhang2021robust, pinto2017robust]. However, the robust RL, trained against bounded perturbations, when exposed to previously unseen threats, their performance often degrades due to lack of adaptability.
I-A3 Meta-Policy Adaptation
Meta-policy approaches generally fall into three families: (1) gradient-based meta-RL, (2) policy ensembling with learned switching among low-level policies, and (3) runtime-assurance architectures that switch from advanced to safe controllers. Gradient-based meta-RL algorithms disentangle task inference from control and adapt via latent context inference, improving sample efficiency over on-policy meta-RL, yet they still require online updates and task-posterior estimation at test time, which can be brittle under distribution shift in the state-space [rakelly2019efficient, gao2023context]. While such meta-policies enable fast adaptation across tasks [rakelly2019efficient], their need for retraining or test-time gradient steps makes them ill-suited for real-time UAV navigation. In offline RL, evaluation-time switching between an RL policy and a behavior-cloning policy using uncertainty estimates has demonstrated adaptation can be done in real-time without retraining [neggatu2025evaluation], and constraint-adaptive policy switching (CAPS) trains a suite of policies with different reward/cost trade-offs and selects among them at test time to satisfy constraints [chemingui2025constraint]. These advances strengthen the case for evaluation-time switching, but they primarily address dataset coverage and safety-constraint satisfaction rather than adversarial OOD threats, and they do not also provide formal robustness guarantees during switching. Runtime-assurance frameworks such as Black-Box Simplex offer supervisory switching for safety [sheikhi2024black], yet they do not directly address cyber threats or value-aware adaptation under adversarial manipulation.
I-B Research Gap and Contribution
I-B1 Research Gap
Prior robust/ensemble RL methods [agarwalcontrastive, osband2016deep] typically average or blend policies to stabilize performance under stochastic variation but do not adapt online to unseen, adversarial shifts. Likewise, adversarially trained approaches [pinto2017robust, mandlekar2017adversarially] presume fixed train-time threat models and lack runtime adaptation. Meta-RL techniques [rakelly2019efficient] enable fast cross-task adaptation but require test-time gradient updates or retraining, which is impractical for real-time UAV operations. Other switching frameworks [neggatu2025evaluation, chemingui2025constraint] demonstrate test-time selection, yet they offer no performance guarantees during switching.
I-B2 Contributions
In this paper, we propose a meta-policy switching framework for UAV deconfliction in adversarial 3D obstacles environment that adaptively switches between pre-trained adversarial action-robust RL policies and a nominal policy. In contrast to existing research areas of adversarial learning, our method uses discounted Thompson sampling (DTS) driven by value-distribution shifts, quantified via the Wasserstein distance, to select among robust policies without any test-time retraining. Unlike the existing meta-policy adaptive cases, here we provide a theoretical analysis showing how our framework minimizes expected regret while yielding adaptive robustness to attacks which causes OOD shifts in observation space, and we establish a novel connection to antifragility, proving hierarchical policy ensemble leads to emergent antifragile behavior. To our knowledge, this is the first integration of DTS with robust RL ensembles for real-time adaptation in UAV systems against self-induced adversarial observations, accompanied by guarantees for adaptive robustness and antifragility.
  • DTS-Guided Meta-Policy Adaptation: We introduce a meta-policy switching framework that, to our knowledge, is the first to couple discounted Thompson sampling (DTS) with ensembles of adversarially robust RL policies. The meta-policy uses self-induced adversarial observations and value-distribution (Wasserstein-1) shift estimates to switch policies online, enabling test-time adaptation without retraining.
  • Theoretical Analysis of Adaptive Robustness and Antifragile Behaviour: We provide a formal regret analysis showing that the DTS sampler asymptotically selects the policy that minimises value-function shifts against self-induced adversarial observations, even under previously unseen perturbations. We further establish conditions under which reduced regret yields antifragile improvement i.e., performance increases as adversarial intensity rises, offering, to our knowledge, the first rigorous characterisation of test-time adaptive robustness and antifragility in adversarial RL.
  • Empirical Validation Under Attacks which causes OOD shifts: Beyond bounded, known attacks, we evaluate against both adversarial attacks which causes low-OOD (projected gradient descent) and high-OOD shift (hijacking-induced GNSS spoofing) in observation space in deconfliction against 3D obstacles. The proposed strategy consistently improves navigation efficiency and conflict-free trajectory rates over robust and vanilla RL baselines, We also performed ablation studies while considering different bandit samplers, and showed that DTS based sampler provided better safety results against the other techniquues.
@@ I-B2 Contributions
Paper Organization:Section II elaborates on UAV deconfliction environment against 3D obstacles. Section III provides a comprehensive examination of action-based RL. In Section IV, the formulation of the robust policies as multi-armed bandits are considered, along with the respective Bernoulli rewards. Section V focuses on the optimality of the Bernoulli rewards with respect to the distribution shift considering the regret minimization problem, and the optimality of adaptation during test-time due to DTS sampler while hence inducing antifragile behaviour. A detailed analysis of the results, including a comparative assessment against robust algorithms based on benchmarks, is provided in Section VI. Finally, Section VII synthesizes the key findings and conclusions.
@@ II-A State Transition Model
The modeling environment is a tactical deconfliction challenge, where the UAV has to navigate from the starting point to the goal while avoiding collision against dynamic 3D obstacles. The resulting trajectory is computed by the interaction between the original flow field due to the target and the interfered flow field due to the obstacle(s). The initial flow field is designed to guide the UAV to the target destination where position vector $\mathbf{P}$, influences the velocity $u\left(\mathbf{P}\right)$ of the UAV as:
@@
where $C$ is the attraction constant towards the goal and $\bar{d}$ being the Euclidean distance between current position $\mathbf{P}$ and the target position $\mathbf{P}_{d}$.
@@
However, the initial flow field is affected due to the presence of dynamic obstacles. The effect due to the motion and shape of the obstacles is encapsulated using $\overline{M}\left(\mathbf{P}\right)$ which is derived in Appendix A. The resulting synthetic disturbance flow field in the dynamic obstacle environment is provided as follows,
@@
The UAV position $\mathbf{P}_{t+1}$ at time $t+1$ is obtained by using the velocity $\overline{u}\left(\mathbf{P}\right)$ from (1) :
@@
The manifestation of the transition model on the UAV is shown in Figure 1. The discrete time UAV position update in (3) becomes the transition model for the agent which is described in the next subsection.
@@ II-B Reinforcement Learning Agent
The subsequent subsections will explain the state, action, and reward for the RL agent used for developing the navigation policy for a given mission for the UAV.
@@ II-B2 Reward
As described in (6), the deconfliction is achieved by prioritizing the shortest path to the target. However, the shortest path may bring the UAV closer to the obstacle surface. To address this risk, a threat zone is defined around the obstacle, where a penalty is imposed if the planned path intersects this zone.
@@ II-B3 Action
The parameters $\left\{\varrho_{0},\varsigma_{0},\vartheta_{n}\right\}$ as defined in Appendix A becomes the action space of the UAV agent.
Remark 1.
While the simulation environment described in this work incorporates complex 3D obstacles, it does not explicitly model several key real-world factors, such as sensor noise, actuator latency, communication packet loss, and environmental disturbances. These aspects can influence policy execution and overall system performance in real UAV deployments. Nonetheless, simulation-based evaluation remains a necessary and widely adopted first step for benchmarking and developing learning algorithms in autonomous systems. It enables controlled, repeatable, and safe assessment prior to real-world deployment. As part of our future work, we intend to extend our evaluation to include these realistic uncertainties and validate the proposed algorithms in hardware-in-the-loop and physical UAV testbeds, thereby aligning with the practical requirements of secure cyber-physical systems
@@ III Action Robust Reinforcement Learning
As illustrated in Figure 2, it is shown to obtain a set of robust policies by varying the parameter $\alpha$. The agent and adversarial policies, denoted as $\pi_{\theta}$ and $\mu_{\omega}$, respectively, generate action $a_{t}$ based on a weighted combination: $\left(1-\alpha_{i}\right)a^{\textup{agent}}$ for the agent and $\alpha a^{\textup{adv}}$ for the adversary. These actions determine the trajectory of UAV navigation in 3D obstacles environment. Upon executing an action $a_{t}$, the agent receives a reward $r_{t}$ while transitioning to a new state $\Phi^{\prime}$. The agent’s overall experience is stored as $e_{t}=\left\{\Phi_{t},a_{t},r_{t},\Phi^{\prime}_{t}\right\}$ in the replay buffer $\mathcal{D}_{\textup{replay}}$. Subsequently, a batch of experiences $\mathcal{D}=\left\{e_{1},\cdots,e_{N}\right\}$ is sampled from the replay buffer to compute the critic and the corresponding actor loss function.
@@
However, to train agent and adversarial policies in a shared action space, a two-player zero-sum Markov game is used [tessler2019action]. Zero-sum Markov game provides a principled framework for training robust agents by explicitly modeling an adversary that aims to minimize the agent’s reward. This formulation captures worst-case interactions between the agent and its environment, enabling the learning of policies that are resilient to strategically optimized disruptions. Unlike methods based on fixed perturbation bounds or stochastic noise, adversarial training in a zero-sum setting simulates dynamic threats based on worst-case vulnerabilities of the agent, thereby anticipating and minimizing potential reward degradation.
In this setting, both players simultaneously select an action and the reward is determined by the current state $\Phi_{t}$ and the joint actions of the agent and adversarial policies, which influences the transition dynamics. The MDP models the game as $\mathcal{M}_{2}=\left(\mathcal{S},\mathcal{A},\mathcal{A}^{\prime},T_{2},\gamma,R_{2},P_{0}\right)$ where $\mathcal{A}$ and $\mathcal{A}^{\prime}$ denote the action spaces available to the agent and the adversary. The state transition probability is given by $T_{2}:\mathcal{S}\times\mathcal{A}\times\mathcal{A}^{\prime}\times\mathcal{S}\rightarrow\mathbb{R}$ and $R_{2}:\mathcal{S}\times\mathcal{A}\times\mathcal{A}^{\prime}\rightarrow\mathbb{R}$ defines the reward for both players. Here, the agent’s policy is represented by $\mu:\mathcal{S}\rightarrow\mathcal{A}$, while the adversary follows the policy $\nu:\mathcal{S}\times\mathcal{A}^{\prime}$ within the environment $\mathcal{M}_{2}$. At each time step $t$, the agent and the adversary select actions $a_{t}=\mu\left(\Phi_{t}\right)$ and ${a_{t}}^{\prime}=\nu\left(\Phi_{t}\right)$, respectively. In the context of a zero-sum game, the agent receives a reward $r_{t}=R_{2}\left(\Phi_{t},a_{t},a^{{}^{\prime}}_{t}\right)$, while the adversary receives a reward of $r_{t}$. In the adversarial game, the following performance objective is considered for optimization.
@@
The objective defined in (11) can be solved using stochastic gradient Langevin dynamics as shown in Appendix D. To determine the number of models in the ensemble set of agent and adversarial policies, entropy is employed as a metric. According to [han2021max, eysenbach2022maximum], the goal to achieve robustness in RL is to minimize the maximum entropy of environmental states. With the increase in adversarial policy’s contribution to the action space, we expect the entropy difference between the exploitation and exploration phases of learning to decrease. However, at higher levels of adversarial contribution, states may exhibit increased entropy, leading to greater randomness in episodic rewards. Let $\mathbf{H}_{\textup{rand}}$ represent the reward entropy during the exploration phase and $\mathbf{H}_{\textup{opt}}$ during the exploitation phase; then, the difference in entropy can be expressed as $\Delta\mathbf{H}=\mathbb{E}\left[\Delta\mathbf{H}^{\textup{rand}}_{\alpha_{i}}-\max\left[\mathbf{H}^{\textup{opt}}_{\alpha_{i}}\right]\right]$. As shown in Figure 2, $\Delta\mathbf{H}$ is used as a metric to find the required number of ensembles, where, if it goes below, $\Delta H_{\textup{th}}$ then we do not consider the given model within the ensemble set. To determine the number of models in the ensemble set of agent and adversarial policies, entropy is employed as a metric. According to [han2021max, eysenbach2022maximum], the goal to achieve robustness in RL is to minimize the maximum entropy of environmental states. With the increase in adversarial policy’s contribution to the action space, we expect the entropy difference between the exploitation and exploration phases of learning to decrease. However, at higher levels of adversarial contribution, states may exhibit increased entropy, leading to greater randomness in episodic rewards. Let $\mathbf{H}_{\textup{rand}}$ represent the reward entropy during the exploration phase and $\mathbf{H}_{\textup{opt}}$ during the exploitation phase; then, the difference in entropy can be expressed as $\Delta\mathbf{H}=\mathbb{E}\left[\Delta\mathbf{H}^{\textup{rand}}_{\alpha_{i}}-\max\left[\mathbf{H}^{\textup{opt}}_{\alpha_{i}}\right]\right]$. As shown in Figure 2, $\Delta\mathbf{H}$ is used as a metric to find the required number of ensembles, where, if it goes below, $\Delta H_{\textup{th}}$ then we do not consider the given model within the ensemble set. The entire algorithm to obtain the ensemble of robust policies is shown in Algorithm 1.